OpenStudy (anonymous):

In 2003, Sammy received $10,000 from her grandmother. Her parents invested all of the money in account that pays interest annually and by 2015, the amount will have grown to $16,900...Write an exponential function to model this growth...Then assume the money continue to grows at the same rate. What would be the balance in 2025?

OpenStudy (anonymous):

Well first how many years are between 2003 and 2015? 12 So x stands for the amount earned per year by leaving the money in the account. 12x + 10,000 = 16,900 -10,000 -10,000 12x = 6,900 ÷ 12 ÷ 12 x = 575 So Sammy gets 575 dollars a year. So to plug that into the second problem it would be; 2025 - 2003 = 22 So, 22 (575) = 10,000 = z 12650 +10,000 = z 22650 = z Am I clear enough?